OpenStudy (anonymous):

@Calculus1 Find the Third Derivative of the function: F(x)=x^2/X^2-1 @Calculus1 Find the Third Derivative of the function: F(x)=x^2/X^2-1 @MIT 18.02 Multiva…

OpenStudy (anonymous):

what is the denomanator. It is confusing maybe put brackets so i can understand it better

OpenStudy (anonymous):

same feeling i do carry as rld x and X what is it

OpenStudy (anonymous):

\[2(x-1)/(x^2-1)^2\]

OpenStudy (anonymous):

thats just y''

OpenStudy (anonymous):

\[f(x)=x^2/x^2-1\] sorry, the capital "X" was a typo. Thank you for the help.

OpenStudy (agreene):

Wait, is it really" \[f(x)=\frac{x^2}{x^2}-1\] or do you mean it is: \[f(x)=\frac{x^2}{x^2-1}\]

OpenStudy (anonymous):

The second of the two.

OpenStudy (agreene):

okay. \[f'(x)= -\frac{2x}{(x^2-1)^2}\] \[f''(x)= \frac{6x^2+2}{(x^2-1)^3}\] \[f'''(x)= -\frac{24x(x^2+1)}{(x^2-1)^4}\] f'(x) can be done with a quotient rule f''(x) can be done by product rule and then chain rule f'''(x) product then chain

OpenStudy (anonymous):

areene you are wrong at f'(x) step

OpenStudy (anonymous):

I got: \[f'(x) = -2x/(2x)^2 \] However, I do not trust my answer.

OpenStudy (anonymous):

i m sorry agree was right

OpenStudy (agreene):

For evaluating f'(x) remember quotient rule: \[\large\frac{d}{dx}\frac{u}{v}=\frac{v\frac{du}{dx}u\frac{dv}{dx}}{v^2}\] take u=x^2 and v = x^2-1 and we arrive at: \[\frac{(x^2-1)(2x)-x^2(\frac{d}{dx}(x^2)-\frac{d}{dx}(1))}{(x^2-1)^2}\] then we move to: \[\frac{(x^2-1)(2x)-x^2(2x)}{(x^2-1)^2}\] combine likes: \[\frac{-2x}{(x^2-1)^2}\]